# A Photorealistic Dataset and Vision-Based Algorithm for Anomaly Detection During Proximity Operations in Lunar Orbit

### v4

@@ V-B ALLO Benchmark Results
The performance of existing learning-based algorithms on the ALLO test set, compared to their MVTec results, is summarized in Table II. Among the learning-based methods, FastFlow achieved the highest P.AP (29.1%), followed by UFlow (20.7%) and STFPM (12.5%). FastFlow also recorded the best I.AUROC (65.9%) and P.AUROC (90.4%) scores on ALLO. However, its performance dropped substantially compared to MVTec, with a 33.5% decrease in I.AUROC and an 8.1% drop in P.AUROC. Overall, normalizing flow-based methods (FastFlow and UFlow) performed best, whereas semi-supervised approaches (DRAEM and DSR) performed poorly.
The results of MRAD and other learning-free algorithms are presented in Table III. MRAD achieved the highest P.AP (62.1%) and I.AUROC (74.9%), outperforming both PAD (P.AP: 57.1%) and the best learning-based method, FastFlow (I.AUROC: 65.9%). Qualitative examples of MRAD, FastFlow, and STFPM inferences on ALLO are shown in Fig. 6.
Ablation studies highlight MRAD’s robustness to parameter variations. Increasing the grid size from its optimized value reduced P.AP by 2.9% and I.AUROC by 7.3%, whereas decreasing the region-growing tolerance caused only a 1.1% P.AP drop with no effect on I.AUROC.
TABLE II: Performance of state-of-the-art anomaly detection algorithms on the ALLO and MVTEC test sets. For each metric, the three top-performing methods are highlighted in red, orange and yellow, respectively.
Dataset Metric DSR [19] Rev. Dist. [21] SuperSimpleNet [39] CFA [22] DRAEM [18] STFPM [25] UFlow [38] FastFlow [24]
MVTec I.AUROC $\uparrow$ 98.2 98.5 98.0 99.5 98.0 95.5 98.9 99.4
P.AUROC $\uparrow$ - 97.8 - 98.5 97.3 97.0 98.7 98.5
P.AP $\uparrow$ 70.2 - - - 68.4 - - -
ALLO I.AUROC $\uparrow$ 55.6 $\pm$ 2.0 57.4 $\pm$ 4.4 49.6 $\pm$ 0.5 51.3 $\pm$ 0.5 55.4 $\pm$ 1.3 61.8 $\pm$ 2.4 60.7 $\pm$ 3.2 65.9 $\pm$ 0.6
P.AUROC $\uparrow$ 69.5 $\pm$ 6.5 75.7 $\pm$ 0.7 79.5 $\pm$ 9.5 84.8 $\pm$ 7.7 77.1 $\pm$ 2.2 87.8 $\pm$ 1.7 85.8 $\pm$ 1.6 90.4 $\pm$ 0.5
P.AP $\uparrow$ 6.9 $\pm$ 0.7 7.0 $\pm$ 2.5 9.3 $\pm$ 1.0 11.0 $\pm$ 2.8 12.1 $\pm$ 1.4 12.5 $\pm$ 1.3 20.7 $\pm$ 1.6 29.1 $\pm$ 1.2
TABLE III: Results of MRAD and other statistical anomaly detection baselines on the ALLO test set. The three top-performing methods are highlighted in red, orange and yellow, respectively.
Algorithm I. AUROC $\uparrow$ P. AUROC $\uparrow$ P. AP $\uparrow$
@@
MRAD (8x8 grids) 75.7 73.6 60.0
MRAD (2x2 grids) 65.6 65.1 59.2
MRAD (small tolerance) 74.9 68.3 61.0
MRAD (large tolerance) 68.5 67.3 54.9
MRAD 74.9 70.0 62.1
@@ V-C1 Limitations of Learning-Based Methods
Furthermore, some methods assume anomaly-free data follow a unimodal distribution [20, 24], which is problematic for datasets where anomalies closely resemble normal data or where normal data is inherently multi-modal [18]. A major limitation of existing methods is their assumption of image consistency. Many were designed for applications with fixed lighting and viewpoints (e.g., industrial defect inspection). As a result, anomaly scores can be unreliable when there are misalignments between training and test images. Normalizing flow-based methods, such as FastFlow [24] and UFlow [38], are better able to handle complex distributions, reducing false positives. However, even these top-performing algorithms struggle to detect anomalies with colors similar to the station.
Learned methods applied to the ALLO dataset struggle to detect anomalies as they attempt to model a global distribution of non-anomalous features. MRAD overcomes this by analyzing each query image individually and leveraging its corresponding anomaly-free reference image.
@@ V-C2 Statistical (Learning-Free) Anomaly Detection
Further analysis of our results demonstrates how a non-learning algorithm can be used for anomaly detection and what information can be leveraged to improve performance. The benefits of using a reference image are observed in the performance increase between RXD and PAD. By calculating background statistics from reference images rather than query images, PAD improved the P.AP of RXD by 39.5%. MRAD further builds on PAD by incorporating noise removal and region growing to reduce false positives and cluster anomalous pixels, leading to a 5% increase in P.AP and an 18.6% improvement in I.AUROC.
The impact of grid size on MRAD’s anomaly detection performance demonstrates how reference images enhance detection. Large grids consider a broader context around a query pixel but often miss fine-grained local anomalies. In contrast, smaller grids better leverage the reference image to detect local deviations. Similarly, low region-growing tolerance values encourage the grouping of nearby anomalous pixels by reinforcing local consistency in anomaly scores. However, high tolerance values tend to incorrectly include non-anomalous pixels, reducing detection accuracy.
Importantly, MRAD’s simple statistical formulation and few tunable parameters make it interpretable and easy to adapt across different operational scenarios. Unlike deep learning approaches, which are often black boxes that can require retraining, MRAD’s explicit use of background statistics provides predictable behavior.
Furthermore, this statistical approach circumvents the significant data dependency of deep learning. Acquiring extensive real data in orbit is prohibitively expensive, and training on synthetic data introduces a sim-to-real gap that can compromise on-orbit performance. MRAD avoids these challenges of data acquisition and domain adaptation by generating its reference image at inference time from a known CAD model and pose information. However, this reliance on a predefined model also presents limitations; because MRAD does not learn from data, its performance is limited by the quality of the reference image and the validity of the underlying statistical assumptions.
@@ VI Conclusion
In this paper, we addressed the task of visual anomaly detection and localization for a space station in lunar orbit, a valuable yet unexplored area for autonomous space operations. We introduced ALLO, the first open-source anomaly detection dataset with synthetic lunar orbit imagery, and established a benchmark for the space domain by evaluating state-of-the-art methods. To overcome the limitations of existing approaches, we proposed MRAD, a novel, statistical algorithm that leverages known camera poses and a CAD model of the station to render an anomaly-free reference image and detect anomalies as statistical deviations. MRAD achieves state-of-the-art performance on the ALLO dataset, outperforming existing methods in both pixel-level AP and image-level AUROC. This work demonstrates the importance of robust anomaly detection and shows how a statistical approach can enhance the autonomy of robotic systems in space.

### v5

@@ V-B ALLO Benchmark Results
The performance of existing learning-based algorithms and MRAD on the ALLO test set is summarized in Table II. Among the learning-based methods, Dinomaly achieves the highest P.AP of 49.4%, followed by FastFlow (29.1%) and UFlow (20.7%), and also attains the best I.AUROC and P.AUROC on ALLO with scores of 78.6% and 91.4%, respectively. However, its performance dropped substantially relative to MVTec, with I.AUROC decreasing by 21.0 points (99.6% to 78.6%) and P.AUROC by 7.0 points (98.4% to 91.4%).
In contrast, MRAD, our learning-free method, outperforms all learning-based and learning-free baselines on ALLO (see Table III), achieving the highest P.AP of 62.9%, best FPR95 of 37.5%, and the second-highest I.AUROC of 75.0%. These results demonstrate that the grid-based dual statistical framework of MRAD enables more accurate detection of fine-grained anomalies. Qualitative examples of FastFlow, STFPM and MRAD are shown in Fig. 7.
TABLE II: Performance of state-of-the-art anomaly detection algorithms on the ALLO test set. For each metric, the three top-performing methods are highlighted in red, orange and yellow, respectively.
Metric DSR [38] Rev. Dist. [9] SuperSimpleNet [29] CFA [22] Dinomaly [14] UniNet [36] STFPM [35] UFlow [33] FastFlow [37] MRAD (ours)
I.AUROC $\uparrow$ 55.6 $\pm$ 2.0 57.4 $\pm$ 4.4 49.6 $\pm$ 0.5 51.3 $\pm$ 0.5 78.6 $\pm$ 0.5 56.8 $\pm$ 2.9 61.8 $\pm$ 2.4 60.7 $\pm$ 3.2 65.9 $\pm$ 0.6 75.0
P.AUROC $\uparrow$ 69.5 $\pm$ 6.5 75.7 $\pm$ 0.7 79.5 $\pm$ 9.5 84.8 $\pm$ 7.7 91.4 $\pm$ 0.1 85.9 $\pm$ 2.2 87.8 $\pm$ 1.7 85.8 $\pm$ 1.6 90.4 $\pm$ 0.5 71.8
[…]
TABLE III: Performance of MRAD with ablations and of non-learning algorithms on the ALLO test set. The optimized settings are a 4 $\mathbin{\!\!\times\!\!}$ 4 grid, a tolerance of 50 MRAD score, kernel size of 20%, and a SD of 5.
Method Ablation I. AUROC $\uparrow$ P. AUROC $\uparrow$ P. AP $\uparrow$
@@
MRAD—Grid Size 8×8 75.7 73.6 60.0
6×6 76.3 71.8 61.3
3×3 72.1 68.1 61.3
2×2 65.6 65.1 59.2
No grids 57.9 58.9 54.7
MRAD—Tolerance 25% 74.9 69.4 60.7
50% 74.9 68.3 61.0
150% 67.5 76.6 58.9
200% 68.5 67.3 54.9
MRAD—Kernel Size 5% 74.9 70.8 61.7
10% 74.9 70.0 62.1
20% 75.0 71.8 62.9
30% 74.9 70.3 60.1
MRAD—Noise SD SD 2 73.0 70.7 62.8
SD 5 75.0 70.6 62.9
SD 10 67.5 77.3 59.4
Optimized 75.0 71.8 62.9
@@ V-C1 Limitations of Learning-Based Methods
In addition, learning-based anomaly detection methods are often limited by their assumptions on the distribution of normal data. Approaches based on unimodality [37] fail when anomalies closely resemble normal samples, or when the normal data distribution is multi-modal. Many algorithms also presume image consistency—fixed lighting and viewpoints typical of industrial inspection tasks. On datasets like ALLO, where lighting and viewing conditions vary and misalignments are common, anomaly scores become unreliable. Feature-reconstruction models such as Dinomaly [14], leverage vision foundation models to learn rich features and achieve the strongest learning-based results on ALLO. Likewise, normalizing flow methods (FastFlow [37], UFlow [33]) better model complex distributions and reduce false positives. Nonetheless, even these advanced methods struggle to detect anomalies that closely resemble the space station, particularly under variable illumination or perspective.
Ultimately, learned methods struggle on ALLO because they attempt to model a single global distribution of non-anomalous features across all views and lighting. In contrast, MRAD analyzes each query image individually via direct reference comparison, eliminating the need for global modeling and enabling robust anomaly localization regardless of lighting or viewing variation.
@@ V-C2 Statistical (Learning-Free) Anomaly Detection
Further analysis of our results demonstrates how a non-learning algorithm can be used for anomaly detection and which components most strongly affect its performance. The value of a reference image is evident in the performance gap between RXD and PAD: computing background statistics from reference rather than query images increases the P.AP of RXD by 39.5%. By comparing RXD scores between the query and reference images, MRAD attenuates the impact of lighting variations and relaxes the Gaussian pixel-distribution assumption of RXD. Combined with noise removal and region growing to suppress false positives and cluster anomalous pixels, MRAD yields a 5.8% increase in P.AP and an 18.7% improvement in I.AUROC over PAD.
The ablation studies identify grid size and region-growing tolerance as the dominant parameters. Increasing the grid size beyond the optimized 4 $\mathbin{\!\!\times\!\!}$ 4 configuration degrades both pixel-level and image-level performance, as large grids favor global context at the expense of fine-grained anomalies. In contrast, smaller grids exploit the reference image more effectively by emphasizing local deviations. Region growing tolerance governs how anomaly scores are aggregated: lower tolerances cause a modest decrease in P.AP with little effect on I.AUROC, while higher tolerances rapidly degrade performance by incorrectly including non-anomalous pixels. For noise handling, kernel size has a larger influence than the Gaussian standard deviation; larger kernels risk removing truly anomalous pixels, whereas smaller kernels fail to suppress background fluctuations. The optimized configuration (a 4 $\mathbin{\!\!\times\!\!}$ 4 grid, tolerance of 50 MRAD score, kernel size of 20%, and standard deviation of 5) provides a practical balance between localization sensitivity and robustness.
These results illustrate how MRAD’s simple statistical formulation and few tunable parameters yield interpretable behavior and straightforward adaptation to different operational scenarios. Unlike deep learning approaches, which typically behave as black boxes and often require retraining, MRAD’s explicit use of background statistics produces predictable responses to changes in illumination and scene content. Moreover, this statistical approach avoids the data dependence of deep learning: acquiring large on-orbit datasets is prohibitively expensive, and training on synthetic imagery introduces a sim-to-real gap that can undermine on-orbit performance. MRAD sidesteps these issues by generating a reference image at inference time from a known CAD model and pose estimate.
This design also introduces limitations. Because MRAD does not learn from data, its performance is constrained by the fidelity of the CAD model, the quality of the rendered reference image, and the validity of its statistical assumptions.
Finally, while the present analysis focuses on free-floating anomalies, since MRAD operates on a per-image basis with a rendered reference, the same framework could naturally be extended to inspection tasks such as detecting structural damage or surface degradation.
@@ VI Conclusion
In this paper, we introduced ALLO, the first open-source anomaly detection dataset with synthetic lunar-orbit imagery, and established a benchmark for space-based anomaly detection by evaluating state-of-the-art methods. We also proposed MRAD, a statistical algorithm that leverages known camera poses and a CAD model to render anomaly-free reference images and detect statistical deviations. On ALLO, MRAD achieves state-of-the-art performance, outperforming existing methods in pixel-level AP, FPR95, and image-level AUROC.